Please help me solve: 1 / 3 ^3 minus [ ( 2 / 3 + 1 / 2 ) divided by 2 / 3 ]
1 / 3 cubed = 1 / 27
2 / 3 + 1 / 2 = 4/6 + 3/6 = 7/6
7/6 divded by 2/3 = 7/6 * 3/2 = 21/12 = 7/4
1/27 - 7/4 = 4/108 - 189/108 = - 185/108 ?????? = 1 and 77/108 ?????

Answer by funmath(2873)

Please help me solve: 1 / 3 ^3 minus [ ( 2 / 3 + 1 / 2 ) divided by 2 / 3 ]
1 / 3 cubed = 1 / 27 (RIGHT!!!!!:-)
2 / 3 + 1 / 2 = 4/6 + 3/6 = 7/6 (RIGHT!!!!!!!!!:-)
7/6 divded by 2/3 = 7/6 * 3/2 = 21/12 = 7/4 (RIGHT!!!!!!:-)
1/27 - 7/4 = 4/108 - 189/108 = - 185/108 (RIGHT!!!!!:-) = 1 and 77/108 ?????
The last thing you did was right, except you forgot the negative -1 and 77/108.
You did a great job---keep it up!!!!
Happy Calculating!!!!